Assume the price of CDs is $16 and the price of DVDs is $22. At those prices, Isabel consumes 19 CDs and 13 DVDs. Her marginal utility from the last CD consumed is 180 and her marginal utility from the last DVD consumed is 212. without changing the combined amount spent on CDs and DVDs, Isabel can increase her utility by consuming ▼ CDs and ▼ DVDs.
Answer #1

Ans) without changing the combined.......... More CDs and less DVDs

Marginal utility of CDs per dollar = 180/16 = 11.25

Marginal utility of DVDs per dollar = 212/22 = 9.64

Since marginal utility of CDs per dollar exceeds marginal utility of DVDs per dollar so more CDs and less DVDs  should be consumed.